OpenStudy (anonymous):

Evaluate:

OpenStudy (anonymous):

\[(d/dx) \int\limits_{x^2}^{10} (dz / (z^2 + 1))\]

OpenStudy (anonymous):

Show steps please

OpenStudy (anonymous):

Well first things first, do you know the antiderivative for the integral?

OpenStudy (anonymous):

Wait lol nvm

OpenStudy (anonymous):

didnt see that part in front.

OpenStudy (anonymous):

\[\Large \int_{x^2}^{10} \frac{dz}{z^2+1}=F(10)-F(x^2) \] Can you tell from here what happens if you differentiate?

OpenStudy (anonymous):

Ah nvm spacelimbus, you win!

OpenStudy (anonymous):

Yeah that makes sense to me the F(10) - F(x^2)

OpenStudy (anonymous):

*smiles* I have had my troubles with these kind of exercises @vf321, hence I had to LaTeX this one out *laughs*

OpenStudy (anonymous):

Differentiate the RHS @ConfusedAboutCalculus, that's what you are looking for.

OpenStudy (anonymous):

So I get the derivative of F(10) - F(x^2)? How?

OpenStudy (anonymous):

Well the first one is a number correct? Any Antiderivative evaluated at a set value: \(F(10)\), if you differentiate you a number, what do you get?

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

exactly, the second one is a bit more tricky, because it's an Antiderivative (which is easy to differentiate) but inside there is another function y=x^2, so you need to apply which rule to differentiate that?

OpenStudy (anonymous):

2x, power rule

OpenStudy (anonymous):

true, but all in all that's the chain rule right?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

good, so do that and you're done

OpenStudy (anonymous):

but what about the dz and z^2 + 1 and all that

OpenStudy (anonymous):

\[ \Large F(10)-F(x^2) \] Differentiate: \[\Large 0 - f(x^2)2x \]

OpenStudy (anonymous):

If your Antiderivative is a function of x, and you differentiate that again, your derivative will be a function of x too.

OpenStudy (anonymous):

I don't understand what you mean. What happens to the Z variables

OpenStudy (anonymous):

Think about an easy example \[ \Large \int_0^x zdz =\frac{x^2}{2} \] If you differentiate the righthandside you get \(x\) which is of the same order than your standard function, just with a different variable.

OpenStudy (anonymous):

in other words, you don't have to worry about the original function anymore it will be the same just with a different substitution for its variables. (Plus the chain rule in this case)

OpenStudy (anonymous):

makes any sense @ConfusedAboutCalculus ?

OpenStudy (anonymous):

Not really.. haha. Since it's the integral from x to 10 of (dz / (z^2 + 1)), I thought you first had to reverse the chain rule and then do F(10) - F(x^2)

OpenStudy (anonymous):

and I thought that f(x) was dz / (z^2 + 1)

OpenStudy (anonymous):

\[ \Large \int_{x^2}^{10} \frac{dz}{z^2+1}=F(10)-F(x^2)\] Do you agree with me that when we differentiate this we get the following (applying Leibniz differential operator on both sides) \[ \large \frac{d}{dx}\int_{x^2}^{10} \frac{dz}{z^2+1}=\frac{d}{dx}(F(10)-F(x^2))\] The LHS is too hard to evaluate directly, fortunately by the fundamental theorem of Calculus it is equal to the RHS of the equation, so lets differentiate that instead. \[\Large \frac{d}{dx}(F(10)- \frac{d}{dx}F(x^2) \]

OpenStudy (anonymous):

so I I thought that if I took the anti derivative of dz / (z^2 + 1), then do F(10) - F(x^2), then find the derivative of that result

OpenStudy (anonymous):

Now for the second one you have to apply the chain rule, derivative of the outside (with respect to the inside) times the derivative of the inside function (inside function is y=x^2, that's the lower bound of your integral)

OpenStudy (anonymous):

So I understand it equals 0 - f(x^2) (2x)

OpenStudy (anonymous):

\[ \Large \frac{d}{dx}F(x^2)=F'(x^2)2x \] and F'(anything)=f(anything)

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

-f(x^2)(2x) ?

OpenStudy (anonymous):

well the result is \[ \Large f(x^2)2x \] And you said already yourself that \[ \Large f(z)= \frac{1}{1+z^2} \] So \[ \Large f(x^2)2x= \frac{2x}{1+x^4}\]

OpenStudy (anonymous):

Back-Substitution.

OpenStudy (anonymous):

why isn't it negative? if it's f(10)(0) - f(x^2)(2x), shouldn't it be -f(x^2)(2x)? then you substitute that?

OpenStudy (anonymous):

of you're right there, it's negative. I did back-substitution but the results wants it negative, as derived above.

OpenStudy (anonymous):

\[ \large 0 - f(x^2)2x=- \frac{2x}{1+x^4}\]

OpenStudy (anonymous):

because x^2 was the lower bound, not the upper bound.

OpenStudy (anonymous):

I think there is an easier way to do this

OpenStudy (anonymous):

If you discovered that, please let me know (-:

OpenStudy (anonymous):

so what If I first did the integral from x^2 to 10 of 1/ (z^2 + 1) and then I got tan^-1(10) - tan^-1(x^2). Then I take the derivative of that and get: 0 - (1/ 1+ x^4)(2x)

OpenStudy (anonymous):

\[\Large \int_{x^2}^0 \frac{1}{z^2+1}dz= \tan^{-1}(z) = - \tan^{-1}(x^2)\]

OpenStudy (anonymous):

final answer is -2x / (1+x^4)

OpenStudy (anonymous):

Very correct, and you know why this works?

OpenStudy (anonymous):

yeah because I used the fundamental theorem of calculus and then took the derivative

OpenStudy (anonymous):

the derivative is over the whole integral

OpenStudy (anonymous):

Simply put, the integral is easy as hell. Try the following \[\Large \frac{d}{dx}\int_0^{x^5} \frac{e^{2z}}{1+z^6}dz \]

OpenStudy (anonymous):

There is no easy integral method for this, but you can still apply the fundamental theorem of Calculus (-:

OpenStudy (anonymous):

I exaggerated a bit with this example, you might want to look for a better one, but as soon as there is an integral you can't easily integrate, you might want to use the FTC instead, over trying to integrate the function manually and then differentiate.

OpenStudy (anonymous):

so the FTC is just a shortcut for integration?

OpenStudy (anonymous):

Well, lets say if you want to discuss the limits of an integral and so on, it's critical points than the FTC of important. Otherwise what you do is very correct and rigorous, but it can't always be applied.

OpenStudy (anonymous):

I see well thank you for your help I learned a lot from the other method you showed me

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

\[\text{Nice}\]